Can Someone please help me, I tried to solve this but I not sure of my answer:

Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?
This is what I got I think it wrong:
3X*X=75
3X^2=75
X^2=75/3
X^2=25
X=SQRT25
X=5 ANSWER FOR THE WIDTH.
3*5=15 ANSWER FOR THE LENGTH.
15*5 = 75
I really need someone help.
Thank you,
J

Answer by checkley77(12844)

3X*X=75
3X^2=75
X^2=75/3
X^2=25
X=SQRT25
X=5 ANSWER FOR THE WIDTH.
3*5=15 ANSWER FOR THE LENGTH.
[Proof:]
15*5 = 75
What is your problem? You have a valid proof.